OpenStudy (anonymous):

The question says to evaluate the integral by making the given substitution. The integral of 5 cos(3x) dx with u=3x

OpenStudy (anonymous):

you mean evaluate the integral by using substitution?

OpenStudy (anonymous):

what would you substitute?

OpenStudy (anonymous):

I don't see where you would substitute....

OpenStudy (anonymous):

ok, how would you integrate cos(3x)?

OpenStudy (anonymous):

The question says to evaluate the integral by making the given substitution. (u=3x)

OpenStudy (anonymous):

\[\int\limits 5\cos(3x) dx, with (u=3x)\]

OpenStudy (anonymous):

you are doing a U substitution \[\int\limits 5\cos(3x) = \int\limits 5\cos(u)\] and your dv=3, so to compensate for the 3 you need a 1/3

OpenStudy (anonymous):

so your integral will look like this \[\int\limits 5/3 \cos(u)\] and take the integral from there and resub back in your u

OpenStudy (anonymous):

sorry forgot the du \[5/3 \int\limits \cos(u) du\]

OpenStudy (anonymous):

-5/3 sin(u) + C = -5/3 sin (3x) + C

OpenStudy (anonymous):

sorry its positive

OpenStudy (anonymous):

How did you get the 5/3 part

OpenStudy (anonymous):

and your dv=3, so to compensate for the 3 you need a 1/3 <-- How did you get that part?